1. The service manager of an appliance store decides to raffle a 50 in. flat screen television. She prepares a list of 500 re show work
Answer #1

2.

3.

Let u denotes the population mean. Hypotheses : To test null hypothesis Ho pe=12 against alternative hypothesis H : > 12 LetThere is sufficient evidence to support the claim that the medium is effective in increasing the bacteria's life expectancy.

4.

Let X denotes the number of times that she hits among 6 times.

Here,

X ~ Binomial(n = 6, p = 3/10 = 0.3)

The probability mass function of X is

P(X = z) = *0.3 * (1 – 0.3)6-1, x = 0,1,2,3,4,5,6

a)

P(X23) = 1 – P(X < 3) = 1 – P(X < 2) = 1 - Σ(3) - * 0.3 * (1 0.3)6-r =1 – 0.74431 = 0.25569

b)

P(X \leq 4) =\sum_{x=0}^{4} \binom{6}{x}*0.3^x*(1-0.3)^{6-x} = 0.989065 \approx 0.9891

c)

P(X =0) =\binom{6}{0}*0.3^0*(1-0.3)^{6-0} = 0.117649 \approx 0.1176

d)

P(X =3) =\binom{6}{3}*0.3^{3}*(1-0.3)^{6-3} = 0.18522